This is my question:
You are creating a collage using two differnet textured papers. One paper cost $.15 per sheet and the other costs $.08 per sheet. You need a total of 12 sheets of paper without spending over $1.50. Let t be the number of sheets of the less expensive paper. Write an inequality that describes the situation. Your inequality should involve only one variable t. Give two different ways you can buy the paper.
Please help.
: This is my question:
You are creating a collage using two differnet textured papers. One paper cost $.15 per sheet and the other costs $.08 per sheet. You need a total of 12 sheets of paper without spending over $1.50. Let t be the number of sheets of the less expensive paper. Write an inequality that describes the situation. Your inequality should involve only one variable t. Give two different ways you can buy the paper.
Please help.

Answer by stanbon(19020)

You are creating a collage using two different textured papers.
One paper cost $.15 per sheet and the other costs $.08 per sheet.
You need a total of 12 sheets of paper without spending over $1.50.
Let t be the number of sheets of the less expensive paper.
--------------------------------------------
Write an inequality that describes the situation. Your inequality should involve only one variable t.
0.08t + 0.15(12-t) <= 1.50
--------------------------------
Give two different ways you can buy the paper.
0.08t + 1.8 - 0.15t <= 1.50
8t + 180 - 15t <= 150
-7t <= -30
t >= 4.2857
4.285 <= t <= 12
-----------------------
Let t = 5 and 12-t = 7
Cost would be 5*8 + 7*15 = 40 + 105 = 145 cents = $1.45
------------------------
Let t= 10 and 12-t = 2
Cost would be 10*8 + 2*15 = 80 + 30 = 110 cents = $1.10
================================
Cheers,
Stan H.